OpenStudy (anonymous):

What number must you add to complete the square? x^2+10x=14

jimthompson5910 (jim_thompson5910):

what's the x coefficient

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

cut that in half to get ______

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

now square that

OpenStudy (anonymous):

25

jimthompson5910 (jim_thompson5910):

that's your answer

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

\[a \div 2 = (\frac{ a }{ 2 })\]\[(\frac{ a }{ 2 })^2 = ?\]